# Improving Instruction-Following in Language Models through Activation Steering

### v1

@@ Abstract
The ability to follow instructions is crucial for numerous real-world applications of language models. In pursuit of deeper insights and more powerful capabilities, we derive instruction-specific vector representations from language models and use them to steer models accordingly. These vectors are computed as the difference in activations between inputs with and without instructions, enabling a modular approach to activation steering. We demonstrate how this method can enhance model adherence to constraints such as output format, length, and word inclusion, providing inference-time control over instruction following. Our experiments across four models demonstrate how we can use the activation vectors to guide models to follow constraints even without explicit instructions and to enhance performance when instructions are present. Additionally, we explore the compositionality of activation steering, successfully applying multiple instructions simultaneously. Finally, we demonstrate that steering vectors computed on instruction-tuned models can transfer to improve base models. Our findings demonstrate that activation steering offers a practical and scalable approach for fine-grained control in language generation.
@@ 1 Introduction
However, user instructions in generative tasks can be more complex and involve multiple parameters that need to be attended to and satisfied during generation (Sun et al., 2023). For example, users may ask for a response to include bullet lists, have a certain number of sentences, or include/exclude specific content. The complexity of such instructions stems from the high number of possible variations, making it impractical to generate post-training data for all scenarios. In the quest for finding more efficient methods for controlling instruction-following and guided by previous research in language model representations, we pose the question: Can we efficiently extract vector representations that encode and control specific instruction-following behavior?
@@
We conduct experiments using the Phi-3 (Abdin et al., 2024), Gemma 2 2B and 9B (Gemma Team, 2024), and Mistral 7B (Jiang et al., 2023) models, focusing on three types of instructions: output format (§ 3), output length (§ 4), and the inclusion/exclusion of specific words (§ 5). Our results on the IFEval dataset (Zhou et al., 2023a) provide evidence that vector representations can encode a wide range of instructions and enhance the model’s instruction-following performance. Notably, we demonstrate that activation steering not only helps models follow constraints when no instruction is provided in the input, but it can also reinforce instruction adherence even when instructions are explicitly present, potentially countering instruction drift (Li et al., 2024). Furthermore, we show that it is possible to simultaneously steer for multiple constraints, such as controlling both format and length (§ 6). Finally, we present surprising evidence that cross-model steering–using vectors computed on an instruction-tuned model to steer a base model– yields better adherence than same-model steering (§ 7), suggesting new possibilities for transferring task-specific skills across models, similar to “task arithmetic” (Ilharco et al., 2023).
@@ 2.2 Steering Procedure
Activation (or representation) engineering involves constructing vectors of activation values that cause desired changes to output text when added to the forward passes of a frozen LLM (Zou et al., 2023). To identify a direction in the model’s residual stream [^2] The model’s residual stream is the per-token hidden state of dimensionality $d_{\mathrm{model}}$ consisting of the sum of all previous component outputs (Elhage et al., 2021). that encodes information about a specific instruction, we use a technique called difference-in-means (Belrose, 2023). This method effectively isolates key feature directions in the model’s internal activations (Marks & Tegmark, 2024; Tigges et al., 2024) and has been used to control various behaviors such as refusal and sycophancy (Arditi et al., 2024; Panickssery et al., 2024). We adapt this method to support finer-grained instructions with multiple possible instantiations (e.g., varying length, varying words to include and exclude). The process involves pairing two versions of the same request: one with only the base query (e.g., “List some facts about Lionel Messi”) and another that additionally includes the instruction we want to represent (e.g., “List some facts about Lionel Messi, ensuring the output is valid JSON”).
@@
This dynamic adjustment allows the model to effectively incorporate the constraint without over- or under-correcting its behavior. For length instructions, which have a more continuous nature, we experiment and show results with different values of $c$, illustrating their impact on the model’s output. Finally, for word-specific constraints, we compute the weight using Eq. 2 and additionally perform a small grid search over neighboring values on a held-out set of examples to fine-tune the steering effect. The steering vector $c\mathbf{u}_{l}$ is then added to the corresponding residual stream layer and the forward pass is resumed with the updated residual stream value $\tilde{\mathbf{x}}^{\prime}_{l}=\mathbf{x}^{\prime}_{l}+c\mathbf{u}_{l}$. This procedure is carried out at a single layer across all token positions, motivated by previous findings that show models tend to deviate from instructions as they generate more tokens (Li et al., 2024). To find the optimal layer for steering, we perform a sweep across a subset of the model’s layers, measuring the effect on a held-out set of queries. Layer and weight selection are further discussed in Apps. D and F.
@@ 2.3 Experimental Setup
Data. We use an augmented version of the IFEval dataset (Zhou et al., 2023a), which consists of 25 distinct instructions, each paired with multiple base queries and expressed in different phrasings, for a total of 541 prompts. To evaluate format instructions, we focus on a subset of 163 examples that specifically relate to the output format. This subset includes instructions such as “The entire output should be wrapped in JSON format,” requests to use a particular language (e.g., “Please respond using only the Punjabi language”), formatting requirements like “Wrap your entire response with double quotation marks,” and casing instructions (e.g., “Answer using lowercase letters only”). A complete list of the instructions used is provided in Appendix A. [^3] Although they are not strictly related to format, we include language constraints (e.g., “the answer should be in German”) in this group, as, like other formatting instructions, they modify the surface-level presentation of the output without affecting the underlying content. For length instructions, we generate prompts by concatenating base queries from IFEval to instructions derived from templates such as “Answer using at most {$n$} sentences.” For word-specific instructions, we use a subset of IFEval containing instructions about the inclusion or exclusion of keywords. The subset has 203 examples, and each example contains a prompt with a single keyword-related instruction. For steering vector computation and layer selection, we construct a separate set of synthetically generated prompts by combining base queries from IFEval with corresponding instruction descriptions to avoid test information leakage. Additional details about the data used are provided in Apps. B and C.
@@
In addition to assessing the model’s instruction-following capabilities, it is important to verify that the model still effectively addresses the base query, even when generating under constraints. To measure the overall quality of the response, we set up a GPT-4o-based evaluation. For each base query $x$, GPT-4o generates a set of five yes/no questions aimed at assessing the quality of the response to $x$. For instance, given the query “Write an essay about the history of Martin Van Buren’s presidency,” GPT-4o generates questions such as “Does the essay provide context on the political, social, and economic climate during Van Buren’s presidency?” and “Is the essay written in clear, grammatically correct English, and does it follow a logical structure?”. Next, given the original query $x$, a model’s response to $x$, and the generated questions we prompt GPT-4o to answer each question, provide a rationale for each yes/no response, or reply with “N/A” if the question is not applicable. We compute the proportion of questions answered positively for a given query and average this score across all queries to obtain a response quality score. We repeat this experiment 3 times and report the mean and standard error. Additional details are provided in Appendix E. [^4] Note that while the model’s response may be influenced by specific instructions, these instructions are not provided to GPT-4o during the evaluation. The goal here is to focus the quality assessment solely on the comprehensiveness and relevance of the response with respect to the base query.
@@ 3 Format Instructions
Table 1: Steering Vector Projections. The projection onto the vocabulary space of the contrastively computed vectors promotes tokens that are semantically related to the respective instruction.
@@
Representation Analysis. To assess whether the model effectively captures instruction-related information in its internal representations, we analyze Phi-3’s residual stream activations at the last token of the input by calculating cosine similarity across three sets of inputs: (1) pairs of inputs that share the same base query, one with and one without the instruction; (2) inputs with different base queries but the same instruction; and (3) inputs with different base queries and no instruction. High similarity in set (2) would suggest that the model captures a shared feature (the instruction), while we expect set (1) to show relatively high similarity due to shared base queries, and set (3) to show lower similarity due to the absence of any shared features. Figure 2 shows how these measures evolve across layers for two instructions. For the “quotation” instruction, where the model is asked to wrap the output in quotation marks, there is a clear difference between sets (2) and (3), indicating that the instruction is partially captured (green vs. red lines in Figure 2a). For the “Urdu language” instruction, set (2) shows higher similarity than set (1), suggesting a strong representation of the instruction (green vs. blue lines in Figure 2b). These results indicate that the model can effectively encode instruction-relevant features at the last input token, which supports steering model behavior based on specific constraints.
@@
Steering Vector Computation. We compute steering vectors for each of the 12 format-related instructions in the IFEval subset and for the 19 language-based instructions specified in the dataset. During the selection of the optimal steering later, we compare the validation score with and without steering to ensure that the steering intervention leads to an improvement. If no layer shows an improvement in the validation score, no steering is applied at test time. Details about the layers selected for steering are provided in Appendix D. To ensure that the steering vectors effectively capture the information related to the instructions, we inspect them by projecting the vectors onto the model’s vocabulary space (nostalgebraist, 2020; Geva et al., 2022). We project the vectors using the model’s unembedding matrix and examine the vocabulary tokens with the highest logit values. § 3 presents the top tokens associated with several of the instructions we consider. We observe that the tokens promoted by the steering vectors are semantically related to the intended instruction, providing an initial validation that the vectors are capturing the desired features.
@@
Steering Results. We first evaluate steering on inputs without explicit text instructions (Figure 3a). In this setting, the instruction-following accuracy without steering hovers around 10% as the input has no information about the instruction. This non-zero accuracy reflects cases where the models incidentally satisfy the instruction, such as when a model rephrases a sentence without using commas, thus accidentally meeting the “no comma” constraint. When steering is applied, we observe a consistent increase in instruction adherence across all models, with accuracy improving to approximately 30%. This shows that the steering vectors encode meaningful information of the instructions and can be effectively used to steer the models toward the intended behavior. Next, we evaluate on inputs with instructions provided in the text (Figure 3b). As expected, the instruction-following accuracy without steering is higher in this setting, ranging between 60% and 90%. Nevertheless, steering still results in a significant performance boost for two out of four models, demonstrating that steering can enhance instruction adherence even when the instructions are explicitly given.
@@
To further analyze which instructions benefit most from steering, we break down Phi-3’s performance by instruction on input without explicit text instructions (Figure 3c). Notably, we observe significant improvements for instructions such as “Lowercase,” which requires the output to be entirely in lowercase characters, and “Highlight Text,” which asks the model to emphasize parts of the response using markdown syntax (e.g., “*{text}*”). However, certain instructions display variability in steering effectiveness. For instance, the “End Checker” instruction requires the model to finish the response with a specific sentence. This ending is often input-dependent; for example, the model might be asked to generate an email ending with “Hope you agree with me,” or to answer a factual question and conclude with “Is there anything else I can help you with?”. Since the steering vector is computed by averaging over all examples containing this instruction, it likely fails to capture the specific sentence required in each individual case, reducing its ability to steer the model toward precisely following this type of instruction.
@@
Response Quality Evaluation. Figure 4 compares the average differences in response quality scores for the four models. The two leftmost groups of bars show the score changes due to steering, with and without input instructions (calculated only for cases where steering is applied). Satisfying specific constraints during generation may lead to less comprehensive responses and hence following instructions is expected to impact the response quality. To show this effect, we also present the differences in quality scores observed when explicit text instructions are added to the base queries without steering. We observe small decreases in quality score due to steering in the setting including instructions (with the exception of Gemma 2B) while the decreases in the setting without instructions are slightly larger but comparable to the effect caused by simply adding the instructions as text. However, we also observe instances where steering compromised the quality of generation–examples where the model generated nonsensical tokens or repeated itself. These failures likely result from suboptimal steering or partially insufficient objectives during the search for layers to intervene, as steering for specific properties can increase perplexity or reduce fluency (Turner et al., 2023; Stickland et al., 2024). We provide some examples from the few cases we observed in Appendix E.
@@ 40 A young girl dreams of a superpower, fights a rival in a dream, and unleashes 1 23
Steering Results. Unlike format instructions, steering the model for length constraints allows for continuous modulation, enabling interpolation between varying degrees of conciseness or verbosity. To explore this, we manually adjust the steering weight $c$ and examine how this affects the response length. On a set of 50 base prompts without any explicit length instructions, we generate responses from Phi-3 using different values of $c$, measuring the distribution of output lengths. As shown in Figure 5a, increasing the value of $c$ effectively shortens the model’s responses: larger steering weights produce increasingly concise outputs. We provide an example of different outputs generated on the same inputs with different steering weights in Table 2. Next, we assess whether steering for length is effective even when the model is provided with explicit length instructions. Using a set of 200 base prompts, we introduce instructions that request the model to limit its response to a maximum of $n\in\{1,\dots,5\}$ sentences. First, we evaluate how often the model adheres to this constraint without steering (light blue bars in Figure 5b). Then, on the same inputs, we apply the steering vector for conciseness with a weight of $c=20$ (dark blue bars in Figure 5b). Across all five values of $n$, we observe a significant and consistent improvement in how often the model’s responses comply with the length constraint. Figure 5c further illustrates how steering shifts the response length distribution toward shorter outputs while still allowing for variability and avoiding overly short responses. Response quality evaluation and results on steering for verbosity are provided in Apps. E and G, respectively.

### v2

@@ Abstract
The ability to follow instructions is crucial for numerous real-world applications of language models. In pursuit of deeper insights and more powerful capabilities, we derive instruction-specific vector representations from language models and use them to steer models accordingly. These vectors are computed as the difference in activations between inputs with and without instructions, enabling a modular approach to activation steering. We demonstrate how this method can enhance model adherence to constraints such as output format, length, and word inclusion, providing inference-time control over instruction following. Our experiments across four models demonstrate how we can use the activation vectors to guide models to follow constraints even without explicit instructions and to enhance performance when instructions are present. Additionally, we explore the compositionality of activation steering, successfully applying multiple instructions simultaneously. Finally, we demonstrate that steering vectors computed on instruction-tuned models can transfer to improve base models. Our findings demonstrate that activation steering offers a practical and scalable approach for fine-grained control in language generation. Our code and data are available at https://github.com/microsoft/llm-steer-instruct.
@@ 1 Introduction
However, user instructions in generative tasks can be more complex and involve multiple parameters that need to be attended to and satisfied during generation (Sun et al., 2023). For example, users may ask for a response to include bullet lists, have a certain number of sentences, or include/exclude specific content. The complexity of such instructions stems from the high number of possible variations, making it impractical to generate post-training data for all scenarios. In the quest for finding more efficient methods for controlling instruction following and guided by previous research in language model representations, we pose the question: Can we efficiently extract vector representations that encode and control specific instruction-following behavior?
@@
We conduct experiments using the Phi-3 (Abdin et al., 2024), Gemma 2 2B and 9B (Gemma Team, 2024), and Mistral 7B (Jiang et al., 2023) models, focusing on three types of instructions: output format (§ 3), output length (§ 4), and the inclusion/exclusion of specific words (§ 5). Our results on the IFEval dataset (Zhou et al., 2023a) provide evidence that vector representations can encode a wide range of instructions and enhance the model’s instruction-following performance. Notably, we demonstrate that activation steering not only helps models follow constraints when no instruction is provided in the input, but it can also reinforce instruction adherence even when instructions are explicitly present, potentially countering instruction drift (Li et al., 2024). Furthermore, we show that it is possible to simultaneously steer for multiple constraints, such as controlling both format and length (§ 6). Finally, we present compelling evidence that cross-model steering–using vectors computed on an instruction-tuned model to steer a base model–is effective and can sometimes yield better adherence than same-model steering (§ 7), suggesting new possibilities for transferring task-specific skills across models, similar to “task arithmetic” (Ilharco et al., 2023).
@@ 2.2 Steering Procedure
Activation (or representation) engineering involves constructing vectors of activation values which cause desired changes to output text when added to the forward passes of a frozen LLM (Zou et al., 2023). To identify a direction in the model’s residual stream [^2] The model’s residual stream is the per-token hidden state of dimensionality $d_{\mathrm{model}}$ consisting of the sum of all previous component outputs (Elhage et al., 2021). that encodes information about a specific instruction, we use a technique called difference-in-means (Belrose, 2023). This method effectively isolates key feature directions in the model’s internal activations (Marks & Tegmark, 2024; Tigges et al., 2024) and has been used to control various behaviors such as refusal and sycophancy (Arditi et al., 2024; Panickssery et al., 2024). We adapt this method to support finer-grained instructions with multiple possible instantiations (e.g., varying length, varying words to include and exclude). The process involves pairing two versions of the same request: one with only the base query (e.g., “List some facts about Lionel Messi”) and another that additionally includes the instruction we want to represent (e.g., “List some facts about Lionel Messi, ensuring the output is valid JSON”).
@@
This dynamic adjustment allows the model to effectively incorporate the constraint without over- or under-correcting its behavior. For length instructions, which have a more continuous nature, we experiment and show results with different values of $c$, illustrating their impact on the model’s output. Finally, for word-specific constraints, we compute the weight using Eq. 2 and additionally perform a small grid search over neighboring values on a held-out set of examples to fine-tune the steering effect. The steering vector $c\mathbf{u}_{l}$ is then added to the corresponding residual stream layer and the forward pass is resumed with the updated residual stream value $\tilde{\mathbf{x}}^{\prime}_{l}=\mathbf{x}^{\prime}_{l}+c\mathbf{u}_{l}$. This procedure is carried out at a single layer across all token positions, motivated by previous findings that show models tend to deviate from instructions as they generate more tokens (Li et al., 2024). To find the optimal layer for steering, we perform a sweep across a subset of the model’s layers, measuring the effect on a held-out set of queries and performing a perplexity-based quality check. Details about this procedure are provided in Appendix E.
@@ 2.3 Experimental Setup
Data. We use an augmented version of the IFEval dataset (Zhou et al., 2023a), which consists of 25 distinct instructions, each paired with multiple base queries and expressed in different phrasings, for a total of 541 prompts. To evaluate format instructions, we focus on a subset of 163 examples that specifically relate to the output format. This subset includes instructions such as “The entire output should be wrapped in JSON format,” requests to use a particular language (e.g., “Please respond using only the Punjabi language”), formatting requirements like “Wrap your entire response with double quotation marks,” and casing instructions (e.g., “Answer using lowercase letters only”). A complete list of the instructions used is provided in Appendix B. [^3] Although they are not strictly related to format, we include language constraints (e.g., “the answer should be in German”) in this group, as, like other formatting instructions, they modify the surface-level presentation of the output without affecting the underlying content. For length instructions, we generate prompts by concatenating base queries from IFEval to instructions derived from templates such as “Answer using at most {$n$} sentences.” For word-specific instructions, we use a subset of IFEval containing instructions about the inclusion or exclusion of keywords. The subset has 203 examples, and each example contains a prompt with a single keyword-related instruction. For steering vector computation and layer selection, we construct a separate set of synthetically generated prompts by combining base queries from IFEval with corresponding instruction descriptions to avoid test information leakage. Additional details about the data used are provided in Apps. C and D.
@@
In addition to assessing the model’s instruction-following capabilities, it is important to verify that the model still effectively addresses the base query, even when generating under constraints. To measure the overall quality of the response, we set up a GPT-4o-based evaluation. For each base query $x$, GPT-4o generates a set of five yes/no questions aimed at assessing the quality of the response to $x$. For instance, given the query “Write an essay about the history of Martin Van Buren’s presidency,” GPT-4o generates questions such as “Does the essay provide context on the political, social, and economic climate during Van Buren’s presidency?” and “Is the essay written in clear, grammatically correct English, and does it follow a logical structure?”. Next, given the original query $x$, a model’s response to $x$, and the generated questions we prompt GPT-4o to answer each question, provide a rationale for each yes/no response, or reply with “N/A” if the question is not applicable. We compute the proportion of questions answered positively for a given query and average this score across all queries to obtain a response quality score. We repeat this experiment 3 times and report the mean and standard error. Additional details are provided in Appendix F. [^4] Note that while the model’s response may be influenced by specific instructions, these instructions are not provided to GPT-4o during the evaluation. The goal here is to focus the quality assessment solely on the comprehensiveness and relevance of the response with respect to the base query.
@@ 3 Format Instructions
Table 1: Steering Vector Projections. The projection onto the vocabulary space of the contrastively-computed vectors promotes tokens that are semantically related to the respective instruction.
@@
Representational Analysis. To assess whether the model effectively captures instruction-related information in its internal representations, we analyze Phi-3’s residual stream activations at the last token of the input by calculating cosine similarity across three sets of inputs: (1) pairs of inputs that share the same base query, one with and one without the instruction; (2) inputs with different base queries but the same instruction; and (3) inputs with different base queries and no instruction. High similarity in set (2) would suggest that the model captures a shared feature (the instruction), while we expect set (1) to show relatively high similarity due to shared base queries, and set (3) to show lower similarity due to the absence of any shared features. Figure 2 shows how these measures evolve across layers for two instructions. For the “quotation” instruction, where the model is asked to wrap the output in quotation marks, there is a clear difference between sets (2) and (3), indicating that the instruction is partially captured (green vs. red lines in Figure 2a). For the “Urdu language” instruction, set (2) shows higher similarity than set (1), suggesting strong representation of the instruction (green vs. blue lines in Figure 2b). These results indicate that the model can effectively encode instruction-relevant features at the last input token, which supports steering model behavior based on specific constraints. Additional representational analyses are provided in Appendix H.
@@
Steering Vector Computation. We compute steering vectors for each of the 12 format-related instructions in the IFEval subset and for the 19 language-based instructions specified in the dataset. During the selection of the optimal steering later, we compare the validation score with and without steering to ensure that the steering intervention leads to an improvement. If no layer shows an improvement in the validation score, no steering is applied at test time. Details about the layers selected for steering are provided in Appendix E. To ensure that the steering vectors effectively capture the information related to the instructions, we inspect them by projecting the vectors onto the model’s vocabulary space (nostalgebraist, 2020; Geva et al., 2022). We project the vectors using the model’s unembedding matrix and examine the vocabulary tokens with the highest logit values. § 3 presents the top tokens associated with several of the instructions we consider. We observe that the tokens promoted by the steering vectors are semantically related to the intended instruction, providing an initial validation that the vectors are capturing the desired features.
@@
Steering Results. We first evaluate steering on inputs without explicit text instructions (Figure 3a). In this setting, the instruction-following accuracy without steering hovers around 10% as the input has no information about the instruction. This non-zero accuracy reflects cases where the models incidentally satisfy the instruction, such as when a model rephrases a sentence without using commas, thus accidentally meeting the “no comma” constraint. When steering is applied, we observe a consistent increase in instruction adherence across all models, with accuracy improving to approximately 30%. This shows that the steering vectors encode meaningful information of the instructions and can be effectively used to steer the models toward the intended behavior. Next, we evaluate on inputs with instructions provided in the text (Figure 3b). As expected, the instruction-following accuracy without steering is higher in this setting, ranging between 60% and 90%. Nevertheless, steering still results in a notceable performance boost for two out of four models, demonstrating that steering can enhance instruction adherence even when the instructions are explicitly given.
@@
To further analyze which instructions benefit most from steering, we break down Phi-3’s performance by instruction on input without explicit text instructions (Figure 3c). Notably, we observe significant improvements for instructions such as “Lowercase,” which requires the output to be entirely in lowercase characters, and “Highlight Text,” which asks the model to emphasize parts of the response using markdown syntax. However, certain instructions display variability in steering effectiveness. For instance, the “End Checker” instruction requires the model to finish the response with a variable, input-dependent sentence (e.g., “Hope you agree with me,” or “Is there anything else I can help you with?”). Because the steering vector is computed by averaging over all examples containing this instruction, it likely fails to capture the specific sentence required in each individual case, reducing its ability to steer the model toward precisely following this type of instruction.
@@
Response Quality Evaluation. Figure 4 compares the average differences in response quality scores for the four models. The two leftmost groups of bars show the score changes due to steering, with and without input instructions (calculated only for cases where steering is applied). Satisfying specific constraints during generation may lead to less comprehensive responses and hence following instructions is expected to impact the response quality. To show this effect, we also present the differences in quality scores observed when explicit text instructions are added to the base queries without steering. We observe small decreases in quality score due to steering in the setting including instructions (with the exception of Gemma 2B) while the decreases in the setting without instructions are slightly larger but comparable to the effect caused by simply adding the instructions as text. However, we also observe instances where steering compromised the quality of generation–examples where the model generated nonsensical tokens or repeated itself. These failures likely result from suboptimal steering or partially insufficient objectives during the search for layers to intervene, as steering for specific properties can increase perplexity or reduce fluency (Turner et al., 2023; Stickland et al., 2024). We provide some examples from the few cases we observed in Appendix F.
@@ 40 A young girl dreams of a superpower, fights a rival in a dream, and unleashes 1 23
Steering Results. Unlike format instructions, steering the model for length constraints allows for continuous modulation, enabling interpolation between varying degrees of conciseness or verbosity. To explore this, we manually adjust the steering weight $c$ and examine how this affects the response length. On a set of 50 base prompts without any explicit length instructions, we generate responses from Phi-3 using different values of $c$, measuring the distribution of output lengths. As shown in Figure 5a, increasing the value of $c$ effectively shortens the model’s responses: larger steering weights produce increasingly concise outputs. We provide an example of different outputs generated on the same inputs with different steering weights in Table 2. Next, we assess whether steering for length is effective even when the model is provided explicit length instructions. Using a set of 200 base prompts, we introduce instructions that request the model to limit its response to a maximum of $n\in\{1,\dots,5\}$ sentences. First, we evaluate how often the model adheres to this constraint without steering (light blue bars in Figure 5b). Then, on the same inputs, we apply the steering vector for conciseness with a weight of $c=20$ (dark blue bars in Figure 5b). Across all five values of $n$, we observe a significant and consistent improvement in how often the model’s responses comply with the length constraint. Figure 5c further illustrates how steering shifts the response length distribution toward shorter outputs while still allowing for variability and avoiding overly short responses. We also conducted experiments with steering for longer outputs and for exact output length, observing positive results (Appendix J). Finally, response quality evaluation results are provided in Appendix F.